What is the range of the shuttlecock launched with an initial speed of 6.3797 m/s at a 58.3699-degree angle with respect to the horizontal from a height of 1.8914 m in a planet where the downward gravitational acceleration has a magnitude of 11.8201 m/s2?


1
Expert's answer
2021-10-06T15:53:42-0400

"l=l_1+l_2,"


"l_1=\frac{v^2\sin2\alpha}g=3.0751~m,"


"h=v\sin\alpha t+\frac{gt^2}2,\implies"

"t=\frac{\sqrt{v^2\sin^2\alpha+2gh}-v\sin\alpha}g=0.2693~s,"

"l_2=v\cos\alpha t=0.9010~m,"


"l=3.9761~m."